Cargando....
|
Mantén pulsado para arrastrar. |
|||
|
Haz clic aquí para cerrar |
|||
Pregunta 1 Informe
Pregunta 2 Informe
Simplify \(2\log \frac{2}{5} - \log \frac{72}{125} + \log 9\)
Pregunta 3 Informe
Pregunta 4 Informe
Pregunta 5 Informe
Evaluate \( \frac{xy^2-x^2y}{x^2-xy^1} \), when x = -2 and y = 2
xy2−x2yx2−xy1
= (−2)(3)2−(−2)2(3)(−2)2−(−2)(3)
= -3
Pregunta 7 Informe
= 1.358
x 1001
= 33.752
= 16.875
= 1678
Pregunta 8 Informe
Find correct to 3 decimal places \(\left(\frac{1}{0.05} + \frac{1}{5.005}\right) - (0.05 x 2.05)\)
500.50.05
- 1025
100.1 - 0.1025
= 99.9975
99.998 to 3 sig. fiq.
Pregunta 9 Informe
What is the product of \( \frac{27}{5^1}(3)^{-3} \) and \( \frac{(1)^{-1}}{5} \)?
Pregunta 10 Informe
If \( \cos x = \sqrt{\frac{a}{b}} \), find coses x
Pregunta 11 Informe
Detalles de la respuesta
Pregunta 12 Informe
Express \(62 + 3\) as decimal correct to 3 significant figures.
Detalles de la respuesta
Pregunta 13 Informe
PMN and PQR are two secants of the circle MQTRN and PT is a tangent. If PM = 5cm, PN = 12cm and PQ = 4.8cm, calculate the respective lengths of PR and PT in centimeters
Pregunta 14 Informe
Pregunta 15 Informe
Simplify \( \frac{3}{x} + \frac{15}{2y} + \frac{6}{xy} \)
Detalles de la respuesta
Pregunta 16 Informe
Pregunta 17 Informe
Pregunta 18 Informe
Evaluate \(x^2(x^2 - 1)^{-\frac{1}{2}} - (x^2 - 1)^{\frac{1}{2}}\)
Pregunta 19 Informe
Pregunta 20 Informe
Pregunta 21 Informe
If w varies inversely as \( \frac{ur}{u+r} \) and is equal to 8 when u = 2 and r = 6, find a relationship between u, v, w.
Pregunta 22 Informe
Rationalize \( \frac{2\sqrt{3}+3\sqrt{2}}{3\sqrt{2}-2\sqrt{3}} \)
Pregunta 23 Informe
In this figure, \(PQ = PR = PS\) and \(\angle SRT = 68^\circ\). Find \(\angle QPS\).
i.e. 2(y + x) + QPS = 360∘
QPS = 360∘ - 2(y + x)
But x + y + 68∘ = 180∘
x + y = 180∘ - 68∘ = 180∘
x + y = 180∘ - 68∘
= 112∘
QPS = 360∘ - 2(112∘ )
360∘ - 224∘ = 136∘
Pregunta 24 Informe
A crate of soft drinks contains 10 bottles of Coca-cola, 8 of Fanta and 6 of sprite. If one bottle is selected at random, what is the probability that it is NOT a Cocacola bottle?
Coca-cola = 10 bottles, Fanta = 8 bottles
Sprite = 6 bottles, Total = 24
P(cola-cola) = 1024
P(not coca-cola) = 1 - 1024
24?1024
= 1424
= 724
Pregunta 25 Informe
In the diagram above, |PQ| = |QR|, |PS| = |RS|, ∠PSR = 30o and ∠PQR = 80o. Find ∠SPQ.
Pregunta 26 Informe
Pregunta 27 Informe
Pregunta 28 Informe
Detalles de la respuesta
Pregunta 29 Informe
Pregunta 30 Informe
Simplify cos2 x (sec2x + sec2 x tan2x)
Pregunta 31 Informe
Simplify \( \frac{\sqrt{1+x}+\sqrt{x}}{\sqrt{1+x}-\sqrt{x}} \)
Detalles de la respuesta
Pregunta 32 Informe
In the figure, PQRS is a square of sides 8cm. What is the area of \( \triangle UVW \)?
Detalles de la respuesta
Pregunta 33 Informe
Pregunta 34 Informe
Pregunta 35 Informe
At what value of x is the function \(y = x^2 - 2x -3\) minimum?
Detalles de la respuesta
Pregunta 36 Informe
Pregunta 37 Informe
Fifty boxes each of 50 bolts were inspected for the number which were defective. The following was the result:
| No. defective per box | 4 | 5 | 6 | 7 | 8 | 9 |
| No. of boxes | 2 | 7 | 17 | 10 | 8 | 6 |
Find the percentage of boxes containing at least 5 defective bolts each.
Pregunta 38 Informe
Simplify \(3\frac{1}{3} - 1\frac{1}{4} x \frac{2}{3} + 1\frac{2}{5}\)
3 - 13
- (54
x 23
) + 125
= 103
- 56
+ 75
= 100−25+4230
= 11730
= 3.9
≈
4
Pregunta 39 Informe
Pregunta 40 Informe
Simplify \( \frac{1}{1+\sqrt{5}} - \frac{1}{1-\sqrt{5}} \)
Pregunta 41 Informe
A car travels from calabar to Enugu, a distance of p km with an average speed of u km per hour and continues to benin, a distance of q km, with an average speed of w km per hour. Find its average speed from Calabar to Benin
Average speed = total DistanceTotal Time
from Calabar to Enugu in time t1, hence
t1 = pu
also from Enugu to Benin
t2 = qw
Average speed = p+qt1+t2
= p+qpu+qw
= p + q × uwpw+qu
= uw(p+q)pw+qu
Pregunta 42 Informe
Detalles de la respuesta
Pregunta 43 Informe
Fifty boxes each of 50 bolts were inspected for the number which were defective. The following was the result:
| No. defective per box | 4 | 5 | 6 | 7 | 8 | 9 |
| No. of boxes | 2 | 7 | 17 | 10 | 8 | 6 |
The mean and the median of the distribution are respectively
Pregunta 44 Informe
Musa borrows N10.00 at 2% per month simple interest and repays N8.00 after 4 months. How much does he still owe?
I = PRT100
= 10×2×4100
= 45
= 0.8
Total amount = N10.80
He pays N8.00
Remainder = 10.80 - 8.00
= N2.80
Pregunta 45 Informe
If the exterior angles of a pentagon are xo, (x + 5)o, (x + 10), (x + 15)o and (x + 20)o, find x
Detalles de la respuesta
Pregunta 46 Informe
Pregunta 47 Informe
Pregunta 48 Informe
¿Te gustaría proceder con esta acción?